# A Grey-box Attack against Latent Diffusion Model-based Image Editing by Posterior Collapse

### v3

@@ Appendix C Ablation Studies and Extra Evaluations
We conduct an ablation study on the variance hyperparameter $v$ in our attack target distribution $p^{\ast}(\bm{z})=\mathcal{N}(\mathbf{0},v\mathbf{I})$. This hyperparameter controls the extent of the latent space manipulation in our attack.
@@
PCA+
For PCA+, our objective is to maintain consistency with the original image while still preventing unauthorized edits. As shown in Table 4, the performance of PCA+ is remarkably similar for both $v=1$ and $v=1\times 10^{-8}$ across most metrics.
@@
1e-8 18.50 18.39 129.09 130.72 0.4233 0.4173 0.5395 0.5401 0.0612 0.0632
Table 4: Ablation study on $v$. In this table, IQAs are taken between the clean image $\bm{x}$ and the adversarial sample after editing $f(\bm{x}_{adv})$. The better IQAs representing the better performance. Notbaly, “clean” for each prompt is the IQA computed between clean image x and edited clean image f (x).
Analyzing the results, we observe that the performance metrics for $v=1$ and $v=1\times 10^{-8}$ are nearly identical in most cases. This similarity suggests that both values of $v$ are capable of achieving the primary goal of PCA+, which is to preserve the original image characteristics while preventing unauthorized edits.
However, a closer examination reveals a subtle but important difference: the FID scores for $v=1$ are consistently slightly lower across all prompts and models. This lower FID indicates that images generated with $v=1$ are marginally closer to the distribution of real images, representing a slightly better preservation of overall image quality and realism.
The marginally superior FID performance of $v=1$ can be attributed to its alignment with the standard normal distribution, which is commonly used as the prior in many diffusion models. This alignment allows for a more natural manipulation of the latent space, resulting in outputs that are subtly more consistent with the expected image distribution.
Based on this observation, we choose $v=1$ for our main experiments with PCA+. While the differences are small, the consistently lower FID scores for $v=1$ suggest a slight advantage in maintaining image realism. This choice allows us to optimally balance our dual objectives of preserving image quality and preventing unauthorized edits, making it the preferred setting for our PCA+ method.
PCA-
As shown in Table 5, setting $v$ close to 0 outperforms $v=1$ across all prompts and both models, as evidenced by all five evaluated metrics. For instance, with prompt P1 on SD14, when v decreases from 1 to 1e-8, PSNR decreases from 18.65 to 18.02, FID increases from 124.33 to 139.11, and SSIM increases from 0.4299 to 0.4441, all indicating enhanced attack effectiveness.
@@
Table 5: Ablation study on $v$. In this table, IQAs are taken between the clean image after editing $f(\bm{x})$ and the adversarial sample after editing $f(\bm{x}_{adv})$. The worse IQAs representing the better performance.
This consistent improvement can be attributed to the tighter concentration of the target distribution around zero when v is smaller. A smaller v encourages the posterior distribution to collapse more severely, leading to more significant disruption in the latent space and, consequently, in the generated images. Our choice of $v=1e-8$ in the main experiments reflects this finding, providing a balance between strong attack performance and numerical stability. This ablation study further demonstrates the flexibility of our method and the importance of carefully tuning hyperparameters for optimal performance.
Ablation on different surrogate VAE
We conducted an ablation study to investigate the impact of using different VAEs as surrogate models in our attack. We compared the performance of VAEs from SD14 and SD15 across five different prompts (P1 to P5) and evaluated their effectiveness on both SD14 and SD15 models. Table 6 presents the results of this study.
[…]
SD14 SD15 SD14 SD15 SD14 SD15 SD14 SD15 SD14 SD15
P1 SD14 18.02 18.02 139.57 136.62 0.4438 0.4454 0.6186 0.6170 0.1105 0.1111
SD15 18.02 18.02 139.11 135.32 0.4441 0.4453 0.6188 0.6171 0.1106 0.1110
P2 SD14 16.74 16.70 72.01 69.71 0.4348 0.4338 0.5998 0.5986 0.1161 0.1169
SD15 16.74 16.70 70.59 69.04 0.4336 0.4329 0.6005 0.5990 0.1156 0.1167
P3 SD14 16.71 16.78 154.99 153.13 0.4226 0.4232 0.6205 0.6206 0.1077 0.1083
SD15 16.71 16.78 156.47 153.85 0.4234 0.4228 0.6202 0.6204 0.1084 0.1079
P4 SD14 17.61 17.49 118.64 115.96 0.3994 0.3931 0.6363 0.6395 0.1051 0.1063
SD15 17.61 17.50 116.48 115.41 0.4004 0.3942 0.6361 0.6394 0.1051 0.1067
P5 SD14 16.81 16.94 134.53 130.34 0.4574 0.4523 0.6032 0.6048 0.1228 0.1205
SD15 16.81 16.95 134.71 129.93 0.4568 0.4524 0.6032 0.6047 0.1227 0.1209
Table 6: Ablation study on different VAEs.
The results show that the choice of surrogate VAE has minimal impact on the attack performance. For most prompts and metrics, the differences between using SD14 VAE and SD15 VAE are negligible. For example, with prompt P1 on SD14, the PSNR values are identical (18.02) for both VAEs, and other metrics show only slight variations.
This consistency across different VAEs suggests that our attack method is robust to the choice of surrogate model. It demonstrates that the vulnerabilities we exploit are inherent to the VAE structure rather than specific to a particular model implementation. This finding has important implications for the generalizability of our attack, as it indicates that an attacker does not need access to the exact VAE used in the target model to mount an effective attack.
The robustness to surrogate model choice further supports the practical applicability of our method, as it can potentially be effective against a wide range of LDM implementations using only publicly available VAE models.
Ablation on $\alpha$
We conducted an ablation study on the learning rate $\alpha$ used in our projected sign gradient ascent. We experimented with $\alpha$ values of 1, 2, and 4 across five different prompts (P1 to P5) and two Stable Diffusion models (SD14 and SD15). Table 7 presents the results of this study.
@@
Table 7: Ablation study on $\alpha$.
As shown in Table 7, $\alpha=2$ generally yields the best performance across most prompts and metrics. For example, with prompt P1 on SD14, when $\alpha=2$, we observe the lowest PSNR (18.02), highest FID (139.11), and highest ACDM (0.1106), indicating the most effective attack. This trend is consistent across different prompts and both models, with a few exceptions where $\alpha=1$ or $\alpha=4$ perform slightly better in specific metrics.
@@
The superior performance of $\alpha=2$ suggests that this learning rate provides an optimal balance between the step size and the number of iterations in our attack algorithm. It allows for sufficiently large updates to effectively perturb the latent space while avoiding overshooting that might occur with larger learning rates.
This ablation study demonstrates the importance of carefully tuning the learning rate in our attack method and justifies our choice of $\alpha=2$ in the main experiments.
Ablation on $t$
@@
To investigate the impact of the number of attack steps on our method’s performance, we conducted an ablation study varying the attack steps $T$. We experimented with $T$ values of 10, 20, and 40 across five different prompts (P1 to P5) and two Stable Diffusion models (SD14 and SD15). Table 8 presents the results of this study.
[…]
SD14 SD15 SD14 SD15 SD14 SD15 SD14 SD15 SD14 SD15
P1 10 19.25 19.26 120.48 117.95 0.4768 0.4781 0.5735 0.5732 0.0732 0.0735
20 18.52 18.52 132.32 129.64 0.4512 0.4525 0.6013 0.6006 0.0941 0.0945
40 18.02 18.02 139.11 135.32 0.4441 0.4453 0.6188 0.6171 0.1106 0.1110
P2 10 17.77 17.74 61.98 61.11 0.4664 0.4657 0.5561 0.5533 0.0812 0.0819
20 17.17 17.12 67.02 65.80 0.4435 0.4421 0.5825 0.5813 0.1005 0.1010
40 16.74 16.70 70.59 69.04 0.4336 0.4329 0.6005 0.5990 0.1156 0.1167
P3 10 17.66 17.70 124.79 123.90 0.4490 0.4498 0.5853 0.5853 0.0759 0.0764
20 17.09 17.14 138.87 140.27 0.4296 0.4292 0.6065 0.6065 0.0938 0.0941
40 16.71 16.78 156.47 153.85 0.4234 0.4228 0.6202 0.6204 0.1084 0.1079
P4 10 18.66 18.55 102.88 100.43 0.4325 0.4273 0.6020 0.6058 0.0700 0.0707
20 18.03 17.91 111.28 108.82 0.4060 0.4010 0.6232 0.6269 0.0896 0.0908
40 17.61 17.50 116.48 115.41 0.4004 0.3942 0.6361 0.6394 0.1051 0.1067
P5 10 17.77 17.94 112.76 111.11 0.4880 0.4850 0.5619 0.5634 0.0904 0.0886
20 17.20 17.36 125.51 122.68 0.4659 0.4616 0.5868 0.5884 0.1083 0.1065
40 16.81 16.95 134.71 129.93 0.4568 0.4524 0.6032 0.6047 0.1227 0.1209
Table 8: Ablation study on $t$.
As shown in Table 8, the performance of our method consistently improves as the number of attack steps increases across all prompts and both models. This trend is evident in all five evaluated metrics: PSNR, FID, SSIM, LPIPS, and ACDM. For example, with prompt P1 on SD14, as t increases from 10 to 40, PSNR decreases from 19.25 to 18.02, FID increases from 120.48 to 139.11, and SSIM decreases from 0.4768 to 0.4441, all indicating enhanced attack effectiveness.
This consistent pattern suggests a strong positive correlation between the number of attack steps and the effectiveness of our Posterior Collapse Attack. The improvement can be attributed to the iterative nature of our approach, allowing for finer-grained adjustments to the adversarial perturbation with each step. Our choice of $T=40$ in the main experiments represents a balance between attack effectiveness and computational efficiency.
Detailed Analysis of VAE Reconstruction Quality
Table 9 presents a comprehensive comparison of IQAs for VAE reconstructions of adversarial samples generated by different methods. The top row ($x^{\text{recon}}$) shows the baseline IQA values for reconstructions of clean images, providing a reference point for the effectiveness of each attack method.
Methods IQAs
[…]
$x^{\text{recon}}$ 29.11 3.73 0.8683 0.1022 0.0286
$x_{\text{adv}}^{\text{recon}}$ AdvDM 24.50 28.02 0.5916 0.4500 0.0384
PG 22.67 63.56 0.5740 0.5550 0.0716
MIST 22.70 62.27 0.5668 0.5545 0.0699
SDS 24.27 25.99 0.6038 0.4322 0.0413
Ours 21.64 72.54 0.5461 0.5838 0.0886
Table 9: Comparison of IQAs for VAE reconstructions of adversarial samples across different methods.
Our proposed Posterior Collapse Attack demonstrates superior performance across all five IQA metrics when considering the reconstructed adversarial samples ($x_{\text{adv}}^{\text{recon}}$). Specifically, our method achieves the lowest PSNR (21.64) and SSIM (0.5461) scores, indicating the most significant degradation in structural similarity between the original and reconstructed adversarial images. The highest FID score (72.54) further corroborates that our method produces reconstructions that deviate substantially from the original image distribution. Moreover, our approach yields the highest LPIPS (0.5838) and ACDM (0.0886) scores, suggesting greater perceptual differences and color discrepancies in the reconstructed adversarial samples. These results consistently outperform other baseline methods.
Our method consistently outperforms other baselines across all IQA metrics, notably without access to the VAE decoder parameters. This constraint makes our approach more realistic and challenging. The significant changes in PSNR, SSIM, FID, LPIPS, and ACDM scores of the reconstructed images demonstrate our method’s effectiveness in inducing substantial distortions in the latent space, which persist through the decoding process. By achieving such disruption without decoder access, our Posterior Collapse Attack presents a practical approach to protecting against unauthorized image manipulation in real-world scenarios where full model access is often unavailable.
Runtime across Different Resolutions
Table 10 presents the runtime and VRAM usage of our proposed method across different image resolutions. These measurements were conducted under the same testing environment as the results presented in Figure 1 of the main paper, using the setup provided by SDS [^2] https://github.com/xavihart/Diff-Protect.
resolution runtime VRAM
256 1.6 3.4G
512 5.9 6.7G
1024 24.7 22.3G
Table 10: Runtime of our proposed method across different resolutions.
All experiments were performed on a single consumer-grade NVIDIA RTX 3090 GPU. Consistent with the main paper, we set the number of iterations t to 40 and the variance v to 1e-8. This consistency in both hardware and hyperparameters eliminates potential variations in runtime due to differences in setup, CUDA versions, or library implementations.
Our method demonstrates efficient performance across various resolutions. For 256x256 images, it requires only 1.6 seconds and 3.4GB of VRAM, scaling up to 24.7 seconds and 22.3GB for 1024x1024 images. The use of an RTX 3090, a common high-end consumer GPU, ensures that these results are representative of real-world performance for many users.
The sub-linear increase in runtime relative to the quadratic increase in pixel count (from 256 to 1024, a 16-fold increase in pixels) indicates good algorithmic efficiency. This performance profile, coupled with the accessibility of the testing hardware, suggests that our approach is practical for a wide range of applications and users, from processing smaller images on modest systems to handling high-resolution images on more powerful setups.
Appendix D Implementation Details
Our proposed method is implemented using PyTorch (Paszke et al. 2019) and the Diffusers library (von Platen et al. 2022). We maintain consistency across all experiments by fixing the random seed to 3407 for PyTorch’s generator (Picard 2021), ensuring reproducibility in all stable diffusion image generations.
Latent Diffusion Models
We utilize two LDMs in our experiments: SD14333https://huggingface.co/CompVis/stable-diffusion-v1-4 and SD15444https://huggingface.co/runwayml/stable-diffusion-v1-5 as victim model. To test the transferability of our method and baselines, we also utilize the SDXL [^5] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 and SD20666stabilityai/stable-diffusion-2. Unless specifically mentioned in the main paper, all inference parameters for these models are set to their default values.
Image Quality Assessment Metrics
We employ five IQA metrics in our evaluation:
  • PSNR: Range set to 1, aligning with PyTorch’s image transformation.
  • FID: Implemented using TorchEval [^7] https://pytorch.org/torcheval/stable/ with default settings.
  • SSIM [^8] https://github.com/Po-Hsun-Su/pytorch-ssim: Gaussian kernel size set to 11.
  • LPIPS [^9] https://github.com/richzhang/PerceptualSimilarity: Utilizing VGG (Simonyan and Zisserman 2015) as the surrogate model.
  • ACDM [^10] https://github.com/ZhongliangGuo/ACDM: Gaussian kernel size set to 11.
Baseline Methods
For comparative analysis, we follow the implementation of SDS2. To ensure fair comparison, we maintain all default parameters except for the attack time step, which is set to 40 across all methods.
Hardware and Environment
The majority of our experiments were conducted on two separate NVIDIA A100-SXM4-80GB GPUs. For transparency and reproducibility, we craft a Docker image based on the diffusers’ official docker [^11] https://hub.docker.com/r/diffusers/diffusers-pytorch-xformers-cuda for the whole project, unless mentioned, all experiments are done in those two machines with the same Docker image.

### v4

@@ VII-A Analysis on $v$
We conduct a analysis on the variance hyperparameter $v$ in our attack target distribution $p^{\ast}(\bm{z})=\mathcal{N}(\mathbf{0},v\mathbf{I})$. This hyperparameter controls the extent of the latent space manipulation in our attack.
@@
VII-A1 PCA+
TABLE VII: Hyperparameters analysis on $v$ for PCA+.
@@ VII-A1 PCA+
1e-8 18.45 18.39 129.09 130.72 0.4033 0.4073 0.5495 0.5501 0.0622 0.0652
For PCA+, our objective is to maintain consistency with the original image while still preventing unauthorized edits. As shown in Tab. VII, the performance of PCA+ is remarkably similar for both $v=1$ and $v=1\times 10^{-8}$ across most metrics. However, the FID scores for $v=1$ are consistently slightly lower across all prompts and models. The marginally superior FID performance of $v=1$ can be attributed to its alignment with the standard normal distribution, which is commonly used as the prior in many diffusion models. Based on this observation, we choose $v=1$ for our main experiments with PCA+.
VII-A2 PCA-
As shown in Tab. VIII, setting $v$ close to 0 outperforms $v=1$ across all prompts and both models, as evidenced by all five evaluated metrics. This consistent improvement can be attributed to the tighter concentration of the target distribution around zero when $v$ is smaller. A smaller $v$ encourages the posterior distribution to collapse more severely, leading to more significant disruption in the latent space and, consequently, in the generated images. Our choice of $v=10^{-8}$ in the main experiments reflects this finding.
TABLE VIII: Hyperparameters analysis on $v$ for PCA-.
@@ VII-A2 PCA-
TABLE IX: Hyperparameters analysis on $\alpha$.
@@
VII-B Analysis on $\alpha$
@@
We conducted a analysis on the learning rate $\alpha$ used in our projected sign gradient ascent. We experimented with $\alpha$ values of 1, 2, and 4 across five different prompts (P1 to P5) and two Stable Diffusion models (SD14 and SD15). Tab. IX presents the results of this study. As shown in Tab. IX, $\alpha=2$ generally yields the best performance across most prompts and metrics. This trend is consistent across different prompts and both models, with a few exceptions where $\alpha=1$ or $\alpha=4$ perform slightly better in specific metrics. The superior performance of $\alpha=2$ suggests that this learning rate provides an optimal balance between the step size and the number of iterations in our attack algorithm. It allows for sufficiently large updates to effectively perturb the latent space while avoiding overshooting that might occur with larger learning rates.
VII-C Analysis on $T$
@@
To investigate the impact of the number of attack steps on our method’s performance, we conducted a analysis varying the attack steps $T$. We experimented with $T$ values of 10, 20, 40, 60, and 80 across five different prompts (P1 to P5) on SD15. Fig. 8 presents the results of this study.
Figure 8: Hyperparameters Analysis on $T$. Performance metrics averaged over five prompts (P1-P5). Error bars indicate standard deviation. All subplots share the same axes: x-axis shows attack steps $T$; y-axis shows IQA scores.
As shown in Fig. 8, the performance of our method improves substantially in early iterations and reaches practical convergence around $T=40$. Specifically, from $T=10$ to $T=40$, we observe significant improvements across all metrics (e.g., PSNR decreases by 1.13, FID increases by approximately 15-20). However, beyond $T=40$, the marginal gains diminish considerably, increasing from $T=40$ to $T=80$ yields less than 5% relative improvement in all metrics, with overlapping error bars indicating statistical insignificance. This convergence behavior demonstrates that $T=40$ provides an effective balance between attack performance and computational efficiency.
VII-D Hyperparameters analysis on $\epsilon$
Figure 9: Hyperparameters analysis on $\epsilon$. (Top) Attack effectiveness reported in mean±std across five prompts. (Bottom) Perturbation cost, measured as the visual quality degradation between original images $\bm{x}$ and adversarial samples $\bm{x}_{adv}$ across five IQA metrics. Each subplot shows how the corresponding metric changes with different $\epsilon$ values (x-axis: $\ell_{\infty}$ budget; y-axis: IQA score).
We evaluate our method across different perturbation budgets ($\epsilon=8,16,32$) to analyze the trade-off between attack effectiveness and perturbation visibility. As shown in Fig. 9, increasing $\epsilon$ from 8 to 16 yields substantial improvement in attack effectiveness with moderate perturbation cost. However, further increasing to $\epsilon=32$ provides minimal additional gain while significantly degrading image quality. This diminishing return pattern validates $\epsilon=16$ as the balanced operating point, which aligns with the standard perturbation budget widely adopted in prior adversarial attack literature [advdm, photoguard, mist, sds].
VIII Conclusion
In this paper, we propose the Posterior Collapse Attack (PCA), a novel protection method against unauthorized LDM-based image editing. PCA requires less than 4% of the model parameters, enabling faster operation and lower VRAM consumption compared to existing approaches. Furthermore, our method demonstrates better transferability across different VAE-based Stable Diffusion architectures.
By extending the study of Posterior Collapse, an architecture-level vulnerability in VAE training, to the inference phase, we identify two distinct phenomena: diffusion collapse and concentration collapse, each serving different protection objectives. We introduce a novel loss function targeting the VAE encoder of LDMs, which can induce different types of collapses in the VAE’s posterior distribution through single hyperparameter adjustment.
Experimental results demonstrate that both variants of PCA significantly outperform previous methods in different protection objective. Our work contributes to the ongoing efforts to secure digital assets in an era of rapidly advancing generative AI, while addressing the socio-technical challenges posed by malicious AI misuse.